The susseptibility and permiability of a diamagnetic substance is

Grade:12th pass

5 Answers

Mahima Kanawat
5 years ago
Dear student... Diamagnetic substances have negative susceptibilities (χ
Mahima Kanawat
5 years ago
Dear student... Diamagnetic substances have negative susceptibilities and Diamagnetic substances have μr less than 1.....WITH REGARDS, MAHIMA (ASKIITIANS FORUM EXPERT)
Khimraj
5 years ago
Diamagnetic substance have following property
  1. -1 \leq Susceptibility
  2. \leq permiability
Hope it clears. 
Khimraj
5 years ago
 
Diamagnetic substance have following property
  1. -1 \leq Susceptibility
  2. \leq permiability
Hope it clears. 
Khimraj
5 years ago
 
Diamagnetic substance have following property
-1 \leq Susceptibility  < 0
\leq permiability   < 1
Hope it clears..................